OpenStudy (anonymous):

Caroline's bedroom is the shape of a rectangular prism that measures 13 feet long, 10 feet wide, and 9 feet high. What is the surface area of her room

OpenStudy (anonymous):

674

OpenStudy (anonymous):

(13*9)*2=234 (13*10)*2=260 (9*10)*2=180

OpenStudy (anonymous):

SA=13*10+9*10+13*9=130+90+117=337 SQ FT

OpenStudy (anonymous):

SORRY THEN 337*2=674SQ FT